# Chapter 0 History of the Observation of Stars

### v3

@@ Abstract
  • The concept of magnitudes as a measure of the brightness of stars was introduced by Hipparchos, too.
@@ 1 Introduction
Most of the physical properties of stars remained uncertain and hidden for a long time. Adopting spectroscopy to astronomical observations turned out to be the key to accessing surface parameters such as temperature, chemical composition, density, and gravitational field strength. Applying these parameters to models, the age of stars can be deduced. Together with 3D positions of stars on a galactic scale, this enables us to study the archaeology of our galaxy.
In the sections of this chapter, we will follow the improvement in observations of stars on a chronological timeline — the History of Observation of Stars — switching between improvements in the determination of different properties of stars.
@@ 2 The brightness of stars in ancient catalogs
Hipparchos’ data of the constellations listed in the above-mentioned commentary on the Phenomena only have rudimentary information about the brightness of stars. The concept of magnitudes, a quantified measure of the brightness of stars, is confirmed for the first time in the poem Astronomica, written by Manilius, a Roman poet of the 1st century AD. Ptolemaios most probably continued to use this system (Cunningham, 2020). In his catalog, he includes mostly numerical data for the brightness of stars, the brightest one starting with 1 mag, and the faintest stars visible with the naked eye with 6 mag, as well as a few descriptive data. As an example, Fig. 2 shows the page from Almagest listing the stars of the constellation Ursa Minor (UMi). Neither a method nor an instrument to determine the stellar magnitudes is mentioned in the Almagest or any other literature of that time. Most probably they used some kind of naked eye estimation.
@@ 2 Magic mechanics improving the accuracy
Tycho Brahe (1546 – 1601) is known for building high-precision instruments. In his early years, Tycho Brahe traveled through Europe, taking studies at many universities and starting astronomical observations. Inadequate observation methods of observatories at the time led to his early involvement with the methodology and instruments for the precise measurement of celestial positions. In April 1575 Tycho Brahe visited the already experienced observer Wilhelm IV and studied his instruments. In the correspondence that followed, Wilhelm IV in turn profited from suggestions for improved measurement accuracy by Brahe, e.g., transversal lines for better reading accuracy. Instead of using clocks [^2] Tycho Brahe did not believe that clocks could have adequate accuracy. He was proved wrong by Jost Bürgi. Brahe designed larger instruments than Wilhelm IV mostly made of brass with high precision scales, the most accurate instrument — mechanically well below one arcmin — was a mural quadrant. Brahe compiled a catalog of 777 stars, which was published in 1602, one year after his death. The full catalog of 1004 stars was published by Kepler in 1627 (Verbunt and van Gent, 2010b). This catalog has an accuracy slightly larger than 2 arcmin.
@@ 1 Telescopes used for positional astronomy
In the 19th century, Friedrich Wilhelm Argelander (1799 - 1875) was able to enhance the positional accuracy to 0.9 arcsec (Eichhorn, 1974) using a meridian circle and carefully correcting precession, nutation, and aberration of the apparent position of stars. The Bonner Durchmusterung, a catalog of 342,198 stars he compiled from his observations, was published between 1859 and 1862.
@@ 3 Further properties of stars emerge
In 1717, Halley compared observations of four stars by Hipparchos and Ptolemaios with those from Brahe and others and despite a warning from Brahe, that the data might be incorrect, concluded that all four stars had shown proper motion.
@@
Determining the brightness of stars remained an unsolved problem until detectors with reproducible sensitivity became available at the end of the 19th century. However, Hipparcos’ magnitude scale was well established for the range of naked–eye vision. It turned out to be a scale of the ratio of intensities, a logarithmic scale. However, the scaling factor was not fixed yet, and not agreed upon between astronomers. With telescope observations, the magnitude scale had to be extended to fainter stars. The typical procedure was to reduce the aperture of the telescope to a size, where a certain star would just not be visible anymore. That aperture size was used as a measure of the magnitude.
In 1856 Norman Pogson (1829 – 1891) reviewed the magnitude scales of various astronomers of his time and in the end came up with a definition of the scale factor that still holds: 5 magnitudes correspond to a brightness ratio of 100 (Pogson, 1856). Still, the absolute calibration of the magnitudes of stars was open.
@@
With the increasing interest in studies of double star systems, the desire for precise measuring devices for very small angular distances arose. The answer was the invention of the heliometer, a double-image micrometer telescope, in the 18th century. It was improved by Fraunhofer in the 1820s. Bessel was impressed by the power of Struve’s heliometer at the Dorpat observatory. In 1829, he put a new heliometer into operation at his observatory in Königsberg and observed the position of the double star system 61 Cyg in 1837 and 1838 relative to two other stars with zero proper motion or parallax. Bessel chose 61 Cyg because it showed a large proper motion which suggested a small distance and thus a large parallax. He was able to find the typical helical motion of 61 Cyg and determined a parallax of about 0.3 arc seconds (Bessel, 1839), the first parallax observation of a star. It’s worth mentioning that Thomas Henderson (1798 – 1844) and Friedrich Georg Otto von Struve (1793 – 1964) were pretty much on a par with Bessel in the ”space race” in the 1830s.
@@ 6 Modern Stellar Astronomy
A major advance in spectroscopy was already used in an early design by Fraunhofer: a diffraction grating spectrometer. Modern spectrometers are stabilized and use very accurate calibration sources, e.g., frequency combs, and reach resolutions powers larger than 100,000 (e.g. HARPS, ESPRESSO). For most of the features interesting to stellar spectroscopy, a resolution power below 100,000 is sufficient. Multi-object spectroscopic telescopes use optical fibers to feed the light of many different sources into a slit spectrometer, used e.g., by SDSS and LAMOST. With a spectral resolution power of 21,000 in the visible range, the 4MOST facility will be able to observe 2436 objects simultaneously. This telescope will start its operation in mid-2024 and will take spectra of thousands of stars within the Milky Way and its neighboring satellite galaxies.
@@ 7 Conclusion
  • Enke et al. (2024) Enke H, Tuvikene T, Groote D, Edelmann H and Heber U (2024). Archives of photographic plates for astronomical use (applause). digitisation of astronomical plates and their integration into the international virtual observatory. Astronomy & Astrophysics .

### v4

@@ Abstract
  • The concept of magnitudes as a quantified measure of the brightness of stars was first mentioned by Manilius.
@@ 1 Introduction
Most of the physical properties of stars remained uncertain and hidden for a long time. Adopting spectroscopy to astronomical observations proved to be the key to accessing surface parameters such as temperature, chemical composition, density, and gravitational field strength. Applying these parameters to models, the age of stars can be deduced. Together with 3D positions of stars on a galactic scale, this enables us to study the archaeology of our galaxy.
In the sections of this chapter, we will follow the improvement in observations of stars on a chronological timeline — the History of the Observation of Stars — switching between improvements in the determination of different properties of stars.
@@ 2 The brightness of stars in ancient catalogs
Hipparchos’ data of the constellations listed in the above-mentioned commentary on the Phenomena only have rudimentary information about the brightness of stars. The concept of magnitudes, a quantified measure of the brightness of stars, is confirmed for the first time in the poem Astronomica, written by Manilius, a Roman poet of the 1st century AD. Ptolemaios most probably continued to use this system (Cunningham, 2020). In his catalog, he includes mostly numerical data for the brightness of stars, the brightest one starting with 1 mag, and the faintest stars visible with the naked eye with 6 mag, as well as a few descriptive data. As an example, Fig. 2 shows the page from Almagest listing the stars of the constellation Ursa Minor (UMi). Neither a method nor an instrument to determine the stellar magnitudes is mentioned in the Almagest or any other literature of that time. Most likely, they used some kind of naked-eye estimation.
@@ 2 Magic mechanics improving the accuracy
Tycho Brahe (1546 – 1601) is known for building high-precision instruments. In his early years, Tycho Brahe traveled through Europe, taking studies at many universities and starting astronomical observations. Inadequate observation methods of observatories at the time led to his early involvement with the methodology and instruments for the precise measurement of celestial positions. In April 1575 Tycho Brahe visited the already experienced observer Wilhelm IV and studied his instruments. In the correspondence that followed, Wilhelm IV in turn profited from suggestions for improved measurement accuracy by Brahe, e.g., transversal lines for better reading accuracy. Instead of using clocks [^2] Tycho Brahe did not believe that clocks could have adequate accuracy. He was proved wrong by Jost Bürgi. Brahe designed larger instruments than Wilhelm IV mostly made of brass with high precision scales; the most accurate instrument — mechanically well below one arcmin — was a mural quadrant. Brahe compiled a catalog of 777 stars, which was published in 1602, one year after his death. The full catalog of 1004 stars was published by Kepler in 1627 (Verbunt and van Gent, 2010b). This catalog has an accuracy slightly larger than 2 arcmin.
@@ 1 Telescopes used for positional astronomy
In the 19th century, Friedrich Wilhelm Argelander (1799 - 1875) was able to further enhance the positional accuracy to 0.9 arcsec (Eichhorn, 1974) using a meridian circle and carefully correcting precession, nutation, and aberration of the apparent position of stars. The Bonner Durchmusterung, a catalog of 342,198 stars he compiled from his observations, was published between 1859 and 1862.
@@ 3 Further properties of stars emerge
In 1717, Halley compared observations of four stars by Hipparchos and Ptolemaios with those from Brahe and others and despite a warning from Brahe, that the data might be incorrect, concluded that all four stars had shown proper motion. But he was wrong.
@@
Determining the brightness of stars remained an unsolved problem until detectors with reproducible sensitivity became available at the end of the 19th century. However, Ptolemaios’ magnitude scale was well established for the range of naked-eye vision. It turned out to be a scale of the ratio of intensities, a logarithmic scale. However, the scaling factor was not fixed yet, and not agreed upon between astronomers. With telescope observations, the magnitude scale had to be extended to fainter stars. The typical procedure was to reduce the aperture of the telescope to a size, where a certain star would just not be visible anymore. That aperture size was used as a measure of the magnitude.
In 1856 Norman Pogson (1829 – 1891) reviewed the magnitude scales of various astronomers of his time and in the end came up with a definition of the scale factor that still holds: 5 magnitudes correspond to a brightness ratio of 100 (Pogson, 1856). Still, the absolute calibration of the magnitudes of stars was an open question at that time. This was solved with the invention of detectors for the brightness of stars.
@@
With the increasing interest in studies of double star systems, the desire for precise measuring devices for very small angular distances arose. The answer was the invention of the heliometer, a double-image micrometer telescope, in the 18th century. It was improved by Fraunhofer in the 1820s. Bessel was impressed by the power of Struve’s heliometer at the Dorpat observatory. In 1829, he put a new heliometer into operation at his observatory in Königsberg and observed the position of the double star system 61 Cyg in 1837 and 1838 relative to two other stars with zero proper motion or parallax. Bessel chose 61 Cyg because it showed a large proper motion which suggested a small distance and thus a large parallax. He was able to find the typical helical motion of 61 Cyg and determined a parallax of about 0.3 arc seconds (Bessel, 1839), the first parallax observation of a star. It’s worth mentioning that Thomas Henderson (1798 – 1844) and Friedrich Georg Otto von Struve (1793 – 1964) were pretty much on a par with Bessel in this ”space race” in the 1830s.
@@ 6 Modern Stellar Astronomy
A major advance in spectroscopy was already used in an early design by Fraunhofer: a diffraction grating spectrometer. Modern spectrometers are stabilized and use very accurate calibration sources, e.g., frequency combs, and reach resolving powers larger than 100,000 (e.g. HARPS, ESPRESSO). For most of the features interesting to stellar spectroscopy, a resolving power below 100,000 is sufficient. Multi-object spectroscopic telescopes use optical fibers to feed the light of many different sources into a slit spectrometer, used e.g., by SDSS and LAMOST. With a spectral resolving power of 21,000 in the visible range, the 4MOST facility will be able to observe 2436 objects simultaneously. This telescope will start its operation in mid-2024 and will take spectra of thousands of stars within the Milky Way and its neighboring satellite galaxies.
@@ 7 Conclusion
  • Enke et al. (2024) Enke H, Tuvikene T, Groote D, Edelmann H and Heber U (2024). Archives of photographic plates for astronomical use (applause). digitisation of astronomical plates and their integration into the international virtual observatory. Astronomy & Astrophysics 687: A165.